-Determine how many solutions exist
-Use either elimination or substitution to find the solutions (if any)
-Graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection Help Thanks :D .
1)4x = 8 and 5y = 15

2) 0.2x + 0.4y = 1.7 and 8.3x - 6.3y = -4.3

Answer by sassylilmommie(4)

4x = 8
5y = 15
4x/4 = 8/4
x = 2
5y = 15
5y/5 = 15/5
y = 3

0.2(6.3) + 0.4(6.3) = 1.7(6.3)
8.3(0.4) – 6.3(0.4) = -4.3(0.4)
1.26x + 2.52y = 10.71
+3.32x – 2.52y = -1.72
4.58x = 8.99

4.58x/4.58 = 8.99/4.58
x = 1.96
0.2(1.96) + 0.4y = 1.7
0.392 + 0.4y - 0.392 = 1.7 – 0.392
0.4y/0.4 = 1.7/0.4
y = 3.25